How long would it take to triple an investment at 4% compounded quarterly
1
Expert's answer
2020-12-13T14:19:36-0500


"A=P(1+\frac{r}{100})^n"


"3P=P(1+\frac{4}{400})^{4n}"

Dividing by P we obtain,

"3=(1.01)^{4n}"

Taking the logarithm of both sides

"log3=4nlog1.01"

This implies that


"4n=\frac{log3}{log1.01}"


"n=\frac{log3}{4log1.01}"

"n=27. 60years"